OpenStudy (anonymous):

Find the length of the curve: y = 1+2x^(3/2)

OpenStudy (anonymous):

\[0 \le x \le 1\]

OpenStudy (anonymous):

L = \[\int\limits_{0}^{1} \sqrt{1 + \frac{ dy }{ dx }}^{2} dx\]

OpenStudy (anonymous):

^yep

OpenStudy (anonymous):

2.268 would be the length when you integrate from 0 to 1

OpenStudy (anonymous):

Well, I to L = \[\int\limits_{0}^{1}\sqrt{1+(3x ^{1/2})}\] And I'm stuck there now

OpenStudy (anonymous):

it'll be sqrt(1+3x)

OpenStudy (anonymous):

So you are integrating from 0 to 1 of sqrt(1+9x). you can get this if you simplify the equation.

OpenStudy (anonymous):

it'll be sqrt(1+9x)

OpenStudy (anonymous):

substitute (1+9x)= t^2

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):

I'm not sure how you get from sqrt(1+3x^(1/2)) to 1+9x

OpenStudy (anonymous):

Its 3x^(3/2)

OpenStudy (anonymous):

actually it's integral of sqrt( 1+ (dy/dx)^2)

OpenStudy (anonymous):

OH I forgot about the square

OpenStudy (anonymous):

and dy/dx= 3 sqrt(x)

OpenStudy (anonymous):

When you square both sides you get the equation that I wrote above

OpenStudy (anonymous):

Okay, I have L = \[\int\limits_{0}^{1} \sqrt{1+9x}dx\]

OpenStudy (anonymous):

yes...now substitute 1+9x= t^2

OpenStudy (anonymous):

or u can write directly the integral of (1+9x)^1/2 = {(1+9x)^(1/2+1)}/9

OpenStudy (anonymous):

if u do the substitution then u'll have to change the limit of integration also.

OpenStudy (anonymous):

so now u can solve by ur own?